Help Save & Exit Check Required Information (The following information applies to the questions displayed below) CommercialSe help would be great!
Answer #1

1 2) 2 Sales 3 Net operating income 4 Margin 5 Turnover 6 Return on investment $ 3,648,000 $ 501,600 13.75% 4.800 66%

A 1 2) 2 Sales =2280000+(2280000*60%) 3 Net operating income =250800+(250800*100%) 4 Margin =+B3/B2 5 Turnover =+B2/760000 6

Return on investment = Net operating income/Average operating assets Return on investment = $501,600/$760,000 Return on inves